OpenStudy (anonymous):

what are two ways you can simplify (sqrt)180 and what is it simplified form.

OpenStudy (lalaly):

determine its prime factors 180=2 x 2 x 3 x 3 x 5\[\sqrt{180}=\sqrt 2 \times \sqrt 2 \times \sqrt 3 \times \sqrt3 \times \sqrt 5\]\[\sqrt {180}=2 \times 3 \sqrt 5 = 6\sqrt 5\]

OpenStudy (anonymous):

thank you :)